Midwest Fabricators Inc. is considering an investment in equipment that will replace direct labor. The equipment has a cost of $143,000 with a $12,000 residual value and a five-year life. The equipment will replace one employee who has an average wage of $51,715 per year. In addition, the equipment will have operating and energy costs of $13,890 per year.

Determine the average rate of return on the equipment, giving effect to straight-line depreciation on the investment. If required, round to the nearest whole percent.
Answer #1

step 1: find depreciation

= cost-salvage/year

=143000-12000 / 5

=26200$

step 2:find average income

savings in cost - depreciation - operating and energy costs

=51715-26200-13890

=11625

step 3 : average investment

= investment at the beginning + investment at the end /2

=143000+12000 / 2

=77500

step 4 : average rate of return

=average annual income/average invesetment

=11625/77500

=15.00%